# A globally convergent gradient method with momentum

### v3

@@ Abstract
In this work, we consider smooth unconstrained optimization problems and we deal with the class of gradient methods with momentum, i.e., descent algorithms where the search direction is defined as a linear combination of the current gradient and the preceding search direction. This family of algorithms includes nonlinear conjugate gradient methods and Polyak’s heavy-ball approach, and is thus of high practical and theoretical interest in large-scale nonlinear optimization. We propose a general framework where the scalars of the linear combination defining the search direction are computed simultaneously by minimizing the approximate quadratic model in the 2 dimensional subspace. This strategy allows us to define a class of gradient methods with momentum enjoying global convergence guarantees and an optimal worst-case complexity bound in the nonconvex setting. Differently than all related works in the literature, the convergence conditions are stated in terms of the Hessian matrix of the bi-dimensional quadratic model. To the best of our knowledge, these results are novel to the literature. Moreover, extensive computational experiments show that the gradient method with momentum here presented is a solid choice to tackle some classes of nonconvex unconstrained problems.
@@ 1 Introduction
Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and are (at least) competitive with a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19]. Moreover, while the conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to still remain superior, the proposed method seems comparable, with some upsides, in the particular, relevant class of multi-layer perceptron training problems [22, 23].
@@ 2 The main idea
  • if $B_{k}=I$, then we obtain the steepest descent direction $-\nabla f(x_{k})$;
  • if $B_{k}=\nabla^{2}f(x_{k})$, then we obtain Newton’s direction $-\left[\nabla^{2}f(x_{k})\right]^{-1}\nabla f(x_{k})$;
@@ 4.1 Conditions on ${B_{k}}$
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of (7).
@@ 4.2 Conditions on $H_{k}$
Then, proposition 3 shows that condition eq. 14 on the eigenvalues of the matrices $B_{k}$ guarantees that the solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ of (8) produces directions $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$ that are gradient-related. However, checking or ensuring that a sequence of $n\times n$ matrices $\{B_{k}\}$ have uniformly bounded eigenvalues can be extremely challenging for large-dimensional optimization problems.
@@
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k},$ where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is the solution of (20).
@@
We note that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$; thus, by using eq. 21 and (25) we have:
$-g_{k}^{T}d_{k}\geq\lambda_{min}(H^{-1}_{k})\biggl{(}\|g_{k}\|^{4}+(g_{k}^{T}s_{k})^{2}\biggr{)},$
@@
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of problem (20). Then, the direction $d_{k}$ satisfies the following conditions:
@@
Recalling eq. 26 and that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$, we have from (30):
$-g_{k}^{T}d_{k}\geq\lambda_{min}(\tilde{H}^{-1})\biggl{(}\|g_{k}\|^{2}+\left(\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\right)^{2}\biggr{)},$
@@ 5 Algorithmic model
    once computed a solution $[\alpha_{k}\ \beta_{k}]^{T}$ of (35), provided one exists, a check on the obtained search direction
@@
7:     if problem (35) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
@@
11:     if gr_dir_found $=$ False then
@@
From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions. Since the Armijo line search is employed within Algorithm 1, the results follow from [29, Proposition 1.2.1] and by [28]. ∎
@@ 7 Computational experiments
In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work. We carried out the tests on two different sets of problems; the results of the two cases are presented in two different subsections.
The code for all the experiments described in this work was written in Python 3.11, exploiting numpy, scipy and pytorch libraries. The software is available at: https://github.com/gliuzzi/GMM.
Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq. 26 holds. As for the line search, we set $\delta=0.5,\;\gamma=10^{-5}$. The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
@@
In a subsequent stage of our experimentation, we also took into account the cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [34, 20, 33, 21]. The cg_descent provides a highly efficient implementation of a conjugate-gradient type method with subspace optimization steps. We used the python wrapper of the original C code of cg_descent v. 6.8, available at https://github.com/alexfikl/pycgdescent, so that we could employ the procedure within the same testing environment as the other considered methods.
@@
7.1 CUTEst problems
We considered a benchmark of 168 problems from the CUTEst test-suite [36]. In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000. In Table 1 those 183 problems are reported along with their dimensions.
Table 1: Collection of 183 unconstrained CUTEst problems ($n$ denotes the number of variables). The problems whose name is crossed out have then been removed from the initial selection as they turned out to be unbounded from below.
@@ 7.1 CUTEst problems
ARGLINA 1000 EDENSCH 2000 OBSTCLBU 9604
ARGLINB 1000 EIGENALS 2550 ODC 4900
ARGLINC 1000 EIGENBLS 2550 ODNAMUR 11130
ARGTRIGLS 1000 EIGENCLS 2652 OSCIGRAD 100000
BDEXP 5000 ENGVAL1 5000 OSCIPATH 5000
BOX 10000 EXTROSNB 1000 PENALTY1 1000
BOXPOWER 10000 FLETBV3M 5000 PENALTY2 1000
BRATU1D 5001 FLETCBV2 5000 PENALTY3 1000
BROWNAL 1000 FLETCBV3 5000 PENTDI 5000
BROYDN3DLS 5000 FLETCHBV 5000 POWELLBC 1000
BROYDNBDLS 5000 FLETCHCR 1000 POWELLSG 5000
CHEBYQAD 1000 FMINSRF2 5625 POWER 10000
CHNRSNBM 1000 FMINSURF 5625 POWERSUM 1000
CLPLATEA 4970 FREUROTH 5000 PRICE3 2000
CURLY30 10000 GENHUMPS 5000 PROBPENL 1000
CVXBQP1 10000 GENROSE 1000 QING 1000
CYCLIC3LS 100002 GENROSEB 1000 QRTQUAD 5000
CYCLOOCFLS 29996 GRIDGENA 5560 QUARTC 5000
CYCLOOCTLS 29996 HILBERTA 1000 RAYBENDL 2046
DEGDIAG 100001 HILBERTB 1000 RAYBENDS 2046
DEGTRID 100001 INDEF 5000 S368 1000
ARWHEAD 5000 INDEFM 100000 SBRYBND 5000
BDQRTIC 5000 INTEQNELS 1002 SCHMVETT 5000
BROYDN7D 5000 JNLBRNG1 9604 SCOSINE 5000
BRYBND 5000 JNLBRNG2 9604 SCURLY10 10000
CHAINWOO 4000 JNLBRNGA 9604 SCURLY20 10000
CLPLATEB 4970 JNLBRNGB 9604 SCURLY30 10000
CLPLATEC 4970 KSSLS 1000 SENSORS 1000
COSINE 10000 LIARWHD 5000 SINEALI 1000
CRAGGLVY 5000 LINVERSE 1999 SINQUAD 5000
CURLY10 10000 LMINSURF 5329 SPARSINE 5000
CURLY20 10000 LUKSAN11LS 1000 SPARSQUR 10000
DEGTRID2 100001 LUKSAN12LS 998 SPIN2LS 1002
DIAGIQB 1000 LUKSAN13LS 998 SPINLS 1327
DIAGIQE 1000 LUKSAN14LS 998 SROSENBR 5000
DIAGIQT 1000 LUKSAN15LS 1000 SSBRYBND 5000
DIAGNQB 1000 LUKSAN16LS 1000 SSC 4900
DIAGNQE 1000 LUKSAN17LS 1000 SSCOSINE 5000
DIAGNQT 1000 LUKSAN21LS 1000 STRTCHDV 1000
DIAGPQB 1000 LUKSAN22LS 1000 TESTQUAD 5000
DIAGPQE 1000 MANCINO 1000 TOINTGSS 5000
DIAGPQT 1000 MCCORMCK 5000 TORSION1 5184
DIXMAANA1 3000 MODBEALE 20000 TORSION2 5184
DIXMAANB 3000 MOREBV 5000 TORSION3 5184
DIXMAANC 3000 NCB20 5010 TORSION4 5184
DIXMAAND 3000 NCB20B 5000 TORSION5 5184
DIXMAANE1 3000 NCVXBQP1 10000 TORSION6 5184
DIXMAANF 3000 NCVXBQP2 10000 TORSIONA 5184
DIXMAANG 3000 NCVXBQP3 10000 TORSIONB 5184
DIXMAANH 3000 NLMSURF 5329 TORSIONC 5184
DIXMAANI1 3000 NOBNDTOR 5184 TORSIOND 5184
DIXMAANJ 3000 NONCVXU2 5000 TORSIONE 5184
DIXMAANK 3000 NONCVXUN 5000 TORSIONF 5184
DIXMAANL 3000 NONDIA 5000 TQUARTIC 5000
DIXMAANM1 3000 NONDQUAR 5000 TRIDIA 5000
DIXMAANN 3000 NONMSQRT 4900 TRIGON1 1000
DIXMAANO 3000 NONSCOMP 5000 TRIGON2 1000
DIXMAANP 3000 OBSTCLAE 9604 VANDANMSLS 1002
DIXON3DQ 10000 OBSTCLAL 9604 VARDIM 1000
DQDRTIC 5000 OBSTCLBL 9604 VAREIGVL 5000
DQRTIC 5000 OBSTCLBM 9604 WOODS 4000
Then, from this set of 183 problems, we removed those problems that are unbounded below, i.e., the problems where at least one solver among the ones considered generates a sequence $\{x_{k}\}$ such that $f(x_{k})\to-\infty$. Those problems are marked in table 1 with a strike-through.
As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
As performance metrics, we considered both the runtime and the number of iterations. In the following, we first report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6. More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively. We also included in this first test scipy conjugate gradient implementation, denoted by CGscipy, as a preliminary baseline.
Figure 1: Performance profiles on all 168 problems.
Figure 2: Performance profiles on the subset of problems where all the solvers find the same best function value.
In fig. 1, we report the performance profiles in terms of time and number of iterations for this comparison comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 168 problems. As we can see, the best performing solver is GMM2 both in terms of time and, together with GMM1 also in terms of iterations. Both solvers GMM2 and GMM3 are also better than CGscipy in terms of time.
Then, in fig. 2, we report the performance profiles considering the subset of problems where each solver finds the same best function value, i.e.
$f^{*}_{i}-f_{L}<10^{-6}\quad\forall\ i\in\cal S,$
where $\cal S$ is the set of solvers. From fig. 2, we note that the performance of GMM methods continues to be clearly superior to that of the baseline conjugate gradient method in terms of computational time, especially in the case of GMM2.
Based on the above results, we will be considering version GMM2 in the rest of the Section as the choice for our GMM algorithm. In what follows, we also allowed a slight degree of non-monotoniticy in GMM, based on the rule proposed in [37].
We then turned to the comparison between our method and some state-of-the-art solvers for unconstrained optimization, namely, L-BFGS [19] and the cg_descent procedure [34, 20, 21]. Performance profiles of solvers GMM and L-BFGS on the entire collection of 168 problems are reported in fig. 3. As we can see, GMM seems to be slightly superior to L-BFGS in both metrics.
Figure 3: Performance profiles for the comparison between GMM and L-BFGS on all the 168 problems.
The picture heavily changes if cg_descent is included in the comparative analysis: as we can see from fig. 5, it appears by far superior than the other solvers considered in the experimentation on this benchmark of problems.
From the one hand, the superior depth and quality of implementation of the cg_descent procedure w.r.t. the simple python code implementing the GMM is probably a component that makes the gap so particularly pronounced; on the other hand, although clearly outperformed, GMM still retains a certain degree of comparability with the state-of-the-art method. This conclusion made us interested in carrying out a further comparison on a different and relevant class of problems.
Figure 4: Performance profiles for the comparison between GMM, cg-descent and L-BFGS on all the 168 problems.
7.2 Multi-layer perceptron training problems
As an additional test for the proposed GMM method we considered a set of problems from supervised learning context, where methods with momentum are particularly popular. In particular, we aim to train multi-layer perceptrons for binary classification tasks [22, 23]. We considered a fully connected networks consisting of two hidden layers of 128 units each. The input layer matches the size of the samples of each dataset. The output layer consists in a single unit with sigmoid activation. The activation function of the hidden units is the hyperbolic tangent.
In total, 30 binary classification tasks are considered. We limit the dimension of the training set by taking at most 10000 samples from each dataset. We do so by randomly drawing the samples from the full dataset, as this allows to maintain the class imbalance of some datasets. A recap of the problems considered is available in table 2.
Table 2: Details of the 30 binary classification problems considered. Value $n$ denotes the number of parameters for the problem.
Dataset $n$ Dataset $n$
a1a 32001 gisette 656769
a2a 32001 ijcnn1 19585
a3a 32385 liver.disorders 17409
a4a 32385 phishing 25473
a5a 32385 sonar 24449
a6a 32385 svmguide1 17281
a7a 32513 svmguide3 19585
a8a 32513 w1a 55169
a9a 32513 w2a 55169
breast.cancer 18049 w3a 55169
cod.rna 17793 w4a 55169
covtype 23681 w5a 55169
diabetes 17793 w6a 55169
fourclass 17025 w7a 55169
german.numer 19841 w8a 55169
The weights in the multi-layer perceptron are initialized using Kaiming uniform initialization [38]. We then ran both GMM and cg_descent from the same starting point, repeating this process with three different random seeds. The objective function is binary cross-entropy. We also added an $\ell_{2}$ regularization term, with a weight $\lambda_{\textbf{reg}}=0.001$. The termination condition is $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$, which is a typical accuracy degree for learning tasks.
We report in fig. 5 the performance profiles in terms of time, number of iterations and value of the objective function obtained for GMM and cg_descent. We considered each pair dataset-seed as a distinct problem in the profiles, so that we have 90 instances overall.
Figure 5: Performance profiles for the comparison between GMM and cg-descent on the multi-layer perceptron training problems.
We can see from these plots that cg_descent is once again superior in terms of runtime; however, we do not observe this time differences in the number of instances solved; moreover, in a quarter of the instances GMM finished faster, and the overall performance gap appears more moderate. Actually, if we look at the number of iterations the two solver really seem comparable - the difference in runtime is arguably due to the higher number of function evaluations required by our GMM.
Another element of interest in these results concerns the final objective function value attained by the two solvers; GMM seems to have a slight advantage in this regard; we find this aspect intriguing, as as the momentum term is often used in machine learning also to escape bad stationary points in the highly nonconvex landscapes resulting from training problems.

### v4

@@ Abstract
In this work, we consider smooth unconstrained optimization problems and we deal with the class of gradient methods with momentum, i.e., descent algorithms where the search direction is defined as a linear combination of the current gradient and the preceding search direction. This family of algorithms includes nonlinear conjugate gradient methods and Polyak’s heavy-ball approach, and is thus of high practical and theoretical interest in large-scale nonlinear optimization. We propose a general framework where the scalars of the linear combination defining the search direction are computed simultaneously by minimizing the approximate quadratic model in the 2 dimensional subspace. This strategy allows us to define a class of gradient methods with momentum enjoying global convergence guarantees and an optimal worst-case complexity bound in the nonconvex setting. Differently than all related works in the literature, the convergence conditions are stated in terms of the Hessian matrix of the bi-dimensional quadratic model. To the best of our knowledge, these results are novel to the literature. Moreover, extensive computational experiments show that the gradient method with momentum here presented is competitive with respect to other popular solvers for nonconvex unconstrained problems.
@@ 1 Introduction
Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19]. Moreover, while the state-of-the-art conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to be clearly superior, the proposed method may represent an interesting baseline for devising a solver competitive with the state of the art.
@@ 2 The main idea
  • if $B_{k}=I$, then we obtain the steepest descent direction $\nabla f(x_{k})$;
  • if $B_{k}=\nabla^{2}f(x_{k})$, then we obtain Newton’s direction $\left[\nabla^{2}f(x_{k})\right]^{-1}\nabla f(x_{k})$;
@@ 4.1 Conditions on $B_{k}$}
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of (7).
@@ 4.2 Conditions on $H_{k}$
Then, proposition 3 shows that condition eq. 14 on the eigenvalues of the matrices $B_{k}$ guarantees that the solution $\alpha_{k}\;\;\beta_{k}]^{T}$ of (8) produces directions $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$ that are gradient-related. However, checking or ensuring that a sequence of $n\times n$ matrices $\{B_{k}\}$ have uniformly bounded eigenvalues can be extremely challenging for large-dimensional optimization problems.
@@
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k},$ where $\alpha_{k}\;\;\beta_{k}]^{T}$ is the solution of (20).
@@
We note that $g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$; thus, by using eq. 21 and (25) we have:
$g_{k}^{T}d_{k}\geq\lambda_{min}(H^{-1}_{k})\biggl(\|g_{k}\|^{4}+(g_{k}^{T}s_{k})^{2}\biggr),$
@@
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of problem (20). Then, the direction $d_{k}$ satisfies the following conditions:
@@
Recalling eq. 26 and that $g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$, we have from (30):
$g_{k}^{T}d_{k}\geq\lambda_{min}(\tilde{H}^{-1})\biggl(\|g_{k}\|^{2}+\left(\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\right)^{2}\biggr),$
@@ 5 Algorithmic model
    once computed a solution $\alpha_{k}\ \beta_{k}]^{T}$ of (35), provided one exists, a check on the obtained search direction
@@
7:  if problem (35) admits solution $\alpha_{k}\;\;\beta_{k}]^{T}$ then
@@
11:  if gr_dir_found == False then
@@
From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions. Since the Armijo line search is employed within Algorithm 1, the results follow from [27, Proposition 1.2.1] and by [26]. ∎
@@ 7 Computational experiments
In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work.
The code for all the experiments described in this work was written in Python 3.11, exploiting numpy and scipy libraries. The software is available at: https://github.com/gliuzzi/GMM.
Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rule discussed in section 6.2, which provided better results than those from section 6.1 and section 6.3 in preliminary experiments; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq. 26 holds. As for the line search, we set $\delta=0.5,\;\gamma=10^{-5}$. We also allowed a slight degree of non-monotoniticy in the line search, based on the rule proposed in [31]. The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
@@
We also took into account the state-of-the-art cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [33, 20, 32, 21]. The cg_descent provides a highly efficient implementation of a conjugate-gradient type method with subspace optimization steps. We used the python wrapper of the original C code of cg_descent v. 6.8, available at https://github.com/alexfikl/pycgdescent, so that we could employ the procedure within the same testing environment as the other considered methods.
@@
We considered a benchmark of 77 unconstrained problems from the CUTEst test-suite [35] with a number of variables $n\geq 1000$. In Table 1 these 77 problems are listed along with their dimension; note that, for problems with a user definable number of variables we always choose the default value.
Table 1: Collection of 77 unconstrained CUTEst problems ($n$ denotes the number of variables).
@@
ARWHEAD 5000 DIXMAANM1 3000 MSQRTBLS 1024
BDQRTIC 5000 DIXMAANN 3000 NCB20 5010
BOX 10000 DIXMAANO 3000 NCB20B 5000
BOXPOWER 20000 DIXMAANP 3000 NONCVXU2 5000
BROYDN3DLS 5000 DIXON3DQ 10000 NONDIA 5000
BROYDN7D 5000 DQDRTIC 5000 NONDQUAR 5000
BROYDNBDLS 5000 DQRTIC 5000 PENALTY1 1000
BRYBND 5000 EDENSCH 2000 POWELLSG 5000
CHAINWOO 4000 EG2 1000 POWER 10000
COSINE 10000 EIGENALS 2550 QUARTC 5000
CRAGGLVY 5000 EIGENBLS 2550 SCHMVETT 5000
CURLY10 10000 EIGENCLS 2652 SINQUAD 5000
CURLY20 10000 ENGVAL1 5000 SINQUAD2 5000
CURLY30 10000 EXTROSNB 1000 SPARSINE 5000
DIXMAANA1 3000 FLETBV3M 5000 SPARSQUR 10000
DIXMAANB 3000 FLETCBV2 5000 SPINLS 1327
DIXMAANC 3000 FLETCHCR 1000 SPMSRTLS 4999
DIXMAAND 3000 FMINSRF2 5625 SROSENBR 5000
DIXMAANE1 3000 FMINSURF 5625 SSBRYBND 5000
DIXMAANF 3000 FREUROTH 5000 SSCOSINE 5000
DIXMAANG 3000 GENHUMPS 5000 TOINTGSS 5000
DIXMAANH 3000 JIMACK 3549 TQUARTIC 5000
DIXMAANI1 3000 KSSLS 1000 TRIDIA 5000
DIXMAANJ 3000 LIARWHD 5000 VAREIGVL 5000
DIXMAANK 3000 MOREBV 5000 WOODS 4000
DIXMAANL 3000 MSQRTALS 1024
As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 1000000, so that we consider a failure each run that ends by hitting this threshold.
As performance metrics, we considered both the runtime and the number of iterations.
The results of our experiments are shown in Figure 1. As we can see, cg_descent procedure appears by far superior than all the other solvers on the considered benchmark of problems. On the other hand, GMM outperforms both L-BFGS and scipy’s CG method.
The superior depth and quality of implementation of the cg_descent procedure w.r.t. the simple python code implementing the GMM is probably a component that makes the gap so particularly pronounced; on the other hand, although clearly outperformed, GMM still retains a certain degree of comparability with the state-of-the-art method. This consideration makes us believe that the integration of preconditioning or subspace minimization steps in GMM might lead to an easily accessible Python code with a comparable efficiency w.r.t. cg_descent.
Figure 1: Performance profiles for the comparison between GMM, cg_descent and the Scipy’s implementations of the conjugate gradient method and L-BFGS, on the 77 problems considered.